One of the chair lifts at a ski resort unloads 1800 skiers per hour at the top of the slope. The ride from the bottom to the top takes 12 minutes. How many skiers are riding on the lift at any given time?
Answer #1

Number of trips in an hour = 60/12 = 5

Chair lift unloading is 1800 skiers per hour

Number of skiers on the lift at any time= 1800/5 = 360